OpenStudy (anonymous):

Loadstar Sensors is a company that makes load/force sensors bases on capacitive sensing technology. The company wants accumulate $30 million for plant expansion 5 years from now. If the company already has $15 million in an investment account for the expansion, how much more must the company add to the account now so that it will have the $30 million 5 years from now? The account earns interest at 10% per year.

OpenStudy (anonymous):

I just need help setting up the problem. Would it be: = 15,000,000 (F/P,10%,5) = 15,000,000(.6209) = 9,313,500

OpenStudy (anonymous):

@jim_thompson5910 @ganeshie8 Familiar with these type of problems? I've been waiting 40 mins for help :C

ganeshie8 (ganeshie8):

compound interest

ganeshie8 (ganeshie8):

we need to solve this :- how much money u need to start wid, if u want it to become $30 million in 5 years, growing at 10% per year

ganeshie8 (ganeshie8):

say the money required is P \(\large 30,000,000 = P(1+.1)^5\)

ganeshie8 (ganeshie8):

solve P first ? can u ?

OpenStudy (anonymous):

P = 30,000,000 / (1+.1)^5

ganeshie8 (ganeshie8):

yup, simplify

OpenStudy (anonymous):

Ahh thanks. I hate these word problems..

ganeshie8 (ganeshie8):

who doesnt :) wat do u get, btw thats not the final answer....

ganeshie8 (ganeshie8):

you need to subtract 15million from watever value u get for P

OpenStudy (anonymous):

18627639.69

OpenStudy (anonymous):

im subtracting because we started off with 15 million right?

ganeshie8 (ganeshie8):

Exactly ! 15 million is already there in th ebank so we just need to add the extra 3milltion+ more

ganeshie8 (ganeshie8):

final answer wud be, 18627639.69 - 15million

OpenStudy (anonymous):

Ok great. So u used the formula. P = F [ 1 / (1+I)^n ]

ganeshie8 (ganeshie8):

yes

OpenStudy (anonymous):

Thanks for the expanation.

OpenStudy (anonymous):

explanation*

ganeshie8 (ganeshie8):

np, you're wlcme :))